# Complexity Scaling Laws for Neural Models using Combinatorial Optimization

### v1

@@ Appendix A Scaling law fit constants
Scaling variable Form Algorithm $n$ $\alpha$ $\gamma$ $k$
Nodes ($n$) $\mu_{\text{cost}}=\left(\frac{n-\gamma}{k}\right)^{\alpha}$ Optimal $0.433$ $-0.215$ $0.906$
Random $1.00$ $0.00$ $1.92$
@@
Optimal $20$ $0.656$ $1.24$ $0.102$
Random $20$ $0.500$ $0.369$ $1.50\times 10^{-2}$
@@ B.3 Complexity scaling assessment
Figure 8: Suboptimality w.r.t. TSP node and spatial dimension scaling evaluated throughout training. For RL experiments, the end of cosine decay is marked with a green vertical line. Subsequent updates use learning rates below $10^{-5}$, which impedes the convergence of models training on larger problem scales. Note that we use smaller 12,800 sample datasets to evaluate each learning curve. Each dataset is distinct since we scale the problem, so small differences between adjacent scales may not be statistically significant. The reduced sample size also (falsely) produces negative suboptimality values for 2D TSP with 5-node and 10-node scales where model performance becomes near-optimal.
@@ Appendix C Loss behavior
For generative modeling, the majority of research on neural scaling laws studies cross-entropy loss [3, 4, 13]. However, in general, we do not expect loss to trend for RL algorithms [14]. Further, the scaling behavior of RL loss is often irrelevant to task performance, whereas cross-entropy loss in language modeling, for example, is directly relevant to task performance. This makes scaling laws for RL loss less interpretable.
@@ C.3 Complexity scaling loss behaviors
One potential explanation is that next-node prediction benefits from the increasing negative supervision: each label for next-node selection provides one positive example and ($n-1$) negative examples. If we were to assume this trend extrapolates, then next-node prediction NLL converges as suboptimality diverges. If so, the joint NLL of full-tour prediction may still diverge given the number node predictions per tour increases, although we did not evaluate this metric. Alternatively, the trend observed for node-level NLL loss may break down shortly after the tested scales. In any case, node-level cross-entropy loss does not correlate with task performance. This complication is avoided when scaling laws are obtained for metrics that correlate with task performance by definition (for example, suboptimality).
@@
Figure 11: Loss w.r.t. TSP node and spatial dimension scaling. Top: Self-supervised critic loss adheres to the corresponding scaling law form observed for suboptimality (Figure 3). Bottom: PPO actor loss only follows a zero-centered trend, as it does for neural scaling experiments (Figure 9). Supervised NLL loss adheres to exponential decay when scaling nodes, deviating from the power growth pattern observed for suboptimality and critic loss. This implies that next-node prediction improves with node scale for SFT models, at least over smaller scales. However, this does not imply that full-tour prediction improves because the number of nodes predicted per tour increases.
@@ D.1.1 Evaluation method
To measure these properties, we perform numerous search descents ($k$) for each TSP problem evaluated. Each descent starts at a random solution and terminates when arriving at a locally optimal solution (w.r.t. one additional 2-opt move). We then count the number of unique local optima found, count the number of search moves required to reach them (basin of attraction size), and evaluate the fraction of descents that arrived at the best-found local optima (relative size of the global optimum’s basin of attraction).
@@ D.1.2 Property convergence as $d\to\infty$
Unlike search depth and BLO visitation rate, the number of local optima has no predictable convergent trend and instead grows non-monotonically. This is not an artifact of limited precision as we observe the same behavior for small-$k$, large-batch evaluations (with a significant underestimation bias). Growth visibly slows by 100 dimensions, and growth is upper bounded by the fixed solution space size, so convergence is plausible but not directly implied by these results.
@@ D.1.3 Property divergence as $n\to\infty$
Figure 13: 2-opt cost landscape properties w.r.t. TSP node and spatial dimension scaling. From top to bottom: 10-node spatial dimension scaling, 20-node spatial dimension scaling, and 2-dimensional node scaling. From left to right: mean local optima count, mean frequency that search arrives at the best-found local optimum (BLO), and mean search depth before arriving at a local optimum. Dimension scaling: The number of local optima exhibits non-monotonic growth that is upper bounded by the fixed size of the solution space. BLO visitation rate and search depth adhere to power decay. We found including a $\gamma$ scale offset improves fit quality, but excluding $\gamma$ still produces reasonable fits. From left to right axis scalings are: [semi-log-x, loglog, loglog]. Node scaling: Due to the expanding solution space, the number of local optima diverges as does the search depth required to reach one. BLO visitation rate decays roughly exponentially. Note that local optima count is underestimated by 100,000 descents per problem [51]. From left to right scalings are: [semi-log-y, semi-log-y, linear].
@@
In this paper, we define 2-exchange as the naive search move that swaps two nodes in the tour sequence. We are not the first to use this definition [59], however the term is sometimes used to refer to 2-opt in informal sources. 2-opt is the distinct search move that inverts a subtour [31].
@@ Appendix E Optimal solution approximation in higher dimensions
[…]
E.1 Approximation method
E.2 Dataset validation
First, we compare surrogate and optimal solution distributions in 2D TSP where we have datasets for the latter via Concorde [17]. Figure 15 shows surrogate solution tour length closely approximates the distribution of optimal solution tour length at both 10 and 20 nodes using best-of-100 and best-of-1000 2-opt descents, respectively. Being normally distributed, we perform two sample t-tests on the means in Table 5. For each node scale, the mean surrogate tour length is slightly below the mean optimal tour length, which is possible since surrogate and optimal datasets sample distinct problems. Thus, along with the surrogate algorithm itself, we are testing whether the smaller size of the surrogate datasets can sufficiently approximate 1.28 million samples from the optimal distribution.
Figure 16: Surrogate solution generation’s estimated mean local optima discovery fraction. Fills show 99% confidence intervals assuming local optima count is normally distributed. True mean local optima count is estimated with small batches of 2-opt local search performing 100,000 descents-per-problem. 10-node surrogate generation never falls below 60% discovery, increasing confidence best-found local optima are usually the true global optima. 20-node generation falls below 10% discovery by 20 spatial dimensions, suggesting best-found local optima at these scales may be suboptimal more often.
@@ E.2 Dataset validation
One problem would arise if the best-found local optimum (BLO) visitation frequency were overestimated by fewer descents in surrogate solution generation. This would imply surrogate solutions require more descents to find the true global optima located in relatively small basins of attraction. Instead, we find the BLO visitation rate of surrogate solution generation is almost identical to that observed in the small batch 100,000 descents-per-problem experiment in Figure 13 (center and top center). The error magnitudes do not exceed 1.5% and 0.7% for 10-node and 20-node dimension scaling arrays, respectively. Alongside the high BLO visitation rates themselves, this suggests global optima are often found with only a few 2-opt descents, a finding consistent with existing analysis of 3-opt [51] (Figure 15 in cited work).
@@
Another problem would arise if surrogate solution generation were to visit only a tiny fraction of local optima. Even with a high BLO visitation rate, eventually the number of missed global optima may accumulate. Figure 16 plots the mean local optima discovery fraction of surrogate solution generation, where we estimate the true mean (divisor) using the 100,000 descents-per-problem experiments. The 10-node surrogate solutions uphold a greater than 60% discovery rate by 100D, increasing confidence in global optima discovery considering the BLO visitation rate maintains values greater than 50%. In contrast, 20-node surrogate solutions’ local optima discovery decreases to around 10% by 100D. This may still be sufficient to find the global optima in the overwhelming majority of problems. But given the 20-node BLO visitation rate sinks to roughly 6% (Figure 13 center), we are of course less confident.
@@
There are three smaller caveats for Figure 16 also worth noting. First, the confidence intervals assume the number of local optima at each problem scale are normally distributed. We did not verify as we only saved summary statistics for these experiments. If the local optima count was (for instance) long-tailed toward larger values, this could add uncertainty to our estimates. The second caveat is we ignore uncertainty in the surrogate solutions’ local optima discovery rate, which is relatively small since the surrogate datasets are orders of magnitude larger than those used to estimate the true discovery rate (2048 and 512 samples for 10 and 20 nodes, respectively). Lastly, 100,000 descents-per-problem may still miss elusive local optima with small basins of attraction. Even so, these local optima are quite unlikely to be global optima if 3-opt trends generalize to higher-dimensional 2-opt solutions [51] (Figure 14 in cited work).
@@ Appendix G Model architecture
Our model design prioritizes simplicity where possible. Besides making experiments easier to implement, this choice was intended to improve the reproducibility of our results and promote their generalization to other deep architectures. All core model components are implemented using PyTorch Transformer modules. Model code, algorithm code, and training scripts are provided in the project repository at https://github.com/lowellw6/complexity-scaling-laws. Specific hyperparameter choices are discussed separately in Appendix H.
@@ G.1 Policy network
We moved forward with these simplifications after obtaining near-optimal performance without them at our experimented problem scales. But Kool et al.’s design has compelling motivation, especially concerning efficiency. Their graph embedding allows each decoder step to attend over the reduced subset of selectable nodes while still referencing a representation of the full problem. This linear ramp down of attention length still incurs $O(n^{3})$ complexity to decode a full tour, as does our approach which maintains constant attention length. [^8] Self-attention has $O(n^{2})$ complexity w.r.t. to the node scale and attention length $n$ [19], which we iterate $n$ times. But pruning previously selected nodes requires only a third the compute all else being equal. Given autoregressive decoding is most of the forward pass, we suggest reconsidering this design choice when experimenting with large node scales.
@@ G.2 Critic network
Attempting to learn more precise credit assignment, we implemented node-level value estimation. We refer to this alternative as compositional value learning, and refer to problem-level value learning as state value learning. We compute compositional values through a straightforward modification to our state-value critic architecture: concatenating the positionally-encoded node selection sequence to the start token for target input. This produces $n+1$ compositional values, each attending over incremental partial tours in one forward pass via a causal mask. The first output remains the problem-level state-value, attending only over the start token and node encodings. The final output is the Q-value, attending over all node selections. We explain how compositional values influence learning objectives in the following subsection.
@@ G.3 PPO using a compositional value baseline
This section describes how we set up the PPO actor-critic loss to learn and leverage node-level compositional value estimates. We use this style of PPO for all RL experiments in the main paper. External to this appendix, we omit these details for simplicity and because we observe no evidence of performance differences between node- and problem-level baselines (Figure 17 green and blue).
@@
In theory, this still allows compositional value targets to diversify in expectation. If rollouts from a given partial tour usually result in low-cost solutions, the expected compositional value target is small, and the inverse is also true. In preliminary experiments, compositional value learning appeared slightly more sample efficient compared to strictly learning problem-level state values, motivating our selection of the former.
However, this distinction disappears when carefully controlling for other variables and averaging over several seeds, at least for 30-node TSP. After initial transient effects, performance distributions are indistinguishable (Figure 17), and we noticed compositional value estimates vary little within a given tour. These findings are not surprising given the critic attempts to learn precise values with a coarse reward signal, an analogue to value estimation in a sparsely rewarded MDP. Further, given tours are sampled on-policy for problems generated online, the critic sees each partial tour exactly once.
@@ Appendix H Hyperparameter settings
Setting HPs effectively by hand would be tedious and rely heavily on intuition. Reinforcement learning often involves more HPs than supervised learning, and setting model HPs for our custom architecture lacks the grounding of existing results. Algorithmic HPO has the added benefit of reducing wall-clock time required to reach sufficient convergence. With finite resources, unbottlenecked compute can only be approximated in the scaling laws that require it. Some of our intuition-informed preliminary experiments required several million gradient updates (and multiple learning rate warm restarts) to sufficiently converge, dramatically slowing research flow. Our HPO-informed experiments converged well in less than half the gradient updates.
@@ H.1 HPO method
We chose BOHB for its balance of anytime and final performance [65]. BOHB is no longer state-of-the-art in final performance but has comparable anytime performance [67] and is relatively easy to implement. We implemented BOHB using Optuna 3.5’s multivariate Tree-structured Parzen Estimator (TPE) [68] for HP sampling (the Bayesian optimization “BO” in BOHB) paired with their Hyberband [69] pruning algorithm (the “HB” in BOHB). Table 7 details the HPs we used for the BOHB algorithm itself along with our evaluation setup. We found BOHB especially sensitive to the minimum gradient updates budgeted per trial. Setting this value too low informs pruning with noise before initial convergence. Setting too high wastes a significant amount of time and compute.
@@
Lastly, we found forcing a warm-start, human-intuition HP configuration for trial 0 allows BOHB to find high-performing regions of the search space considerably faster than when using only random startup trials.

### v2

@@ 20 $n$
Scaling variable Form Algorithm $n$ $\alpha$ $\gamma\quad$ $k$
Nodes ($n$) $\mu_{\text{cost}}=\left(\frac{n-\gamma}{k}\right)^{\alpha}$ Optimal $0.433$ 0.215-0.215 $0.906$
Random $1.000$ $0.000$ $1.918$
@@
Optimal 2020 $0.656$ $1.24$ $0.102$
Random 2020 $0.500$ $0.369$ $1.50\times 10^{-2}$
@@ B.3 Complexity scaling assessment
Figure 8: Suboptimality w.r.t. TSP node ($n$) and spatial dimension ($d$) scaling evaluated throughout training. For RL experiments, the end of cosine decay is marked with a green vertical line. Subsequent updates use learning rates below $10^{-5}$, which impedes the convergence of models training on larger problem scales. Note that we use smaller 12,800 sample datasets to evaluate each learning curve. Each dataset is distinct since we scale the problem, so small differences between adjacent scales may not be statistically significant. The reduced sample size also (falsely) produces negative suboptimality values for 2D TSP with 5-node and 10-node scales where model performance becomes near-optimal.
@@ Appendix C Loss behavior
For generative modeling, the majority of research on neural scaling laws studies cross-entropy loss [3, 4, 13]. However, in general, we do not expect loss to trend for RL algorithms [14]. Further, the scaling behavior of RL loss is often irrelevant to task performance, whereas cross-entropy loss in language modeling, for example, is directly relevant to task performance. This detachment from performance makes scaling laws for RL loss less interpretable.
@@ C.3 Complexity scaling loss behaviors
One potential explanation is that next-node prediction benefits from the increasing negative supervision: each label for next-node selection provides one positive example and (n1n-1) negative examples. If we were to assume this trend extrapolates, then next-node prediction NLL converges as suboptimality diverges. If so, the joint NLL of full-tour prediction may still diverge given the number node predictions per tour increases, although we did not evaluate this metric. Alternatively, the trend observed for node-level NLL loss may break down shortly after the tested scales. In any case, node-level cross-entropy loss does not correlate with task performance. This complication is avoided when scaling laws are obtained for metrics that correlate with task performance by definition (for example, suboptimality).
Figure 11: Loss w.r.t. TSP node and spatial dimension scaling. Top: Self-supervised critic loss adheres to the corresponding scaling law form observed for suboptimality (Figure 3). Bottom: PPO actor loss only follows a zero-centered trend, as it does for neural scaling experiments (Figure 9). Supervised NLL loss adheres to exponential decay when scaling nodes, deviating from the power growth pattern observed for suboptimality and critic loss. This trend implies that next-node prediction improves with node scale for SFT models, at least over smaller scales. However, this trend does not imply that full-tour prediction improves because the number of nodes predicted per tour increases.
@@ D.1.1 Evaluation method
To measure these properties, we perform numerous search descents ($k$) for each TSP problem evaluated. Each descent starts at a random tour and terminates when arriving at a locally optimal tour (w.r.t. one additional 2-opt move). We then count the number of unique local optima found, count the number of search moves required to reach them (basin of attraction size), and evaluate the fraction of descents that arrived at the best-found local optima (relative size of the global optimum’s basin of attraction).
@@ D.1.2 Property convergence as $d\to\infty$
Unlike search depth and BLO visitation rate, the number of local optima has no predictable convergent trend and instead grows non-monotonically. This is not an artifact of sample size as we observe the same behavior for small-$k$, large-batch evaluations (with a significant underestimation bias). Growth visibly slows by 100 dimensions, and growth is upper bounded by the fixed solution space size, so convergence is plausible but not directly implied by these results.
Figure 13: 2-opt cost landscape properties w.r.t. TSP node and spatial dimension scaling. From top to bottom: 10-node spatial dimension scaling, 20-node spatial dimension scaling, and 2-dimensional node scaling. From left to right: mean local optima count, mean frequency that search arrives at the best-found local optimum (BLO), and mean search depth before arriving at a local optimum. Dimension scaling: The number of local optima exhibits non-monotonic growth that is upper bounded by the fixed size of the solution space. BLO visitation rate and search depth adhere to power decay. We found that including a $\gamma$ scale offset improves fit quality, but excluding $\gamma$ still produces reasonable fits. From left to right axis scalings are: [semi-log-x, log-log, log-log]. Node scaling: Due to the expanding solution space, the number of local optima diverges as does the search depth required to reach one. BLO visitation rate decays roughly exponentially. Note that local optima count is underestimated by 100,000 descents per problem [52]. From left to right scalings are: [semi-log-y, semi-log-y, linear].
@@ D.1.3 Property divergence as $n\to\infty$
In this paper, we define "2-exchange" as the naive search move that swaps two nodes in the tour sequence. We are not the first to use this definition [66]; however, the term is sometimes used to refer to 2-opt in informal sources. 2-opt is the distinct search move that inverts a subtour [31].
@@ E.1 Approximation method
E.2 Dataset validation
First, we compare surrogate and optimal solution distributions in 2D TSP where we have datasets for the latter via Concorde [17]. Figure 15 shows that surrogate solution tour length closely approximates the distribution of optimal solution tour length at both 10 and 20 nodes using best-of-100 and best-of-1000 2-opt descents, respectively. Being normally distributed, we perform two sample t-tests on the means in Table 6. For each node scale, the mean surrogate tour length is slightly below the mean optimal tour length, which is possible since surrogate and optimal datasets sample distinct problems. Thus, along with the surrogate algorithm itself, we are testing whether the smaller size of the surrogate datasets can sufficiently approximate 1.28 million samples from the optimal distribution.
@@ E.2 Dataset validation
$\mathbb{P}(t<\text{t-value})$ 0.377 0.183
$\mathbb{P}(t>\text{t-value})$ 0.623 0.817
@@
One problem would arise if the best-found local optimum (BLO) visitation frequency were overestimated by fewer descents in surrogate solution generation. This overestimation would imply surrogate solutions require more descents to find the true global optima located in relatively small basins of attraction. Instead, we find the BLO visitation rate of surrogate solution generation is almost identical to that observed in the small batch 100,000 descents-per-problem experiment in Figure 13 (center and top center). The error magnitudes do not exceed 1.5% and 0.7% for 10-node and 20-node dimension scaling arrays, respectively. Alongside the high BLO visitation rates themselves, this suggests global optima are often found with only a few 2-opt descents, a finding consistent with existing analysis of 3-opt [52] (Figure 15 in cited work).
@@
Another problem would arise if surrogate solution generation were to visit only a tiny fraction of local optima. Even with a high BLO visitation rate, eventually the number of missed global optima may accumulate. Figure 16 plots the mean local optima discovery fraction of surrogate solution generation, where we estimate the true mean (divisor) using the 100,000 descents-per-problem experiments. The 10-node surrogate solutions uphold a greater than 60% discovery rate by 100 dimensions, increasing confidence in global optima discovery considering the BLO visitation rate maintains values greater than 50%. In contrast, 20-node surrogate solutions’ local optima discovery decreases to around 10% by 100 dimensions. This discovery rate may still be sufficient to find the global optima in the overwhelming majority of problems. But given the 20-node BLO visitation rate sinks to roughly 6% (Figure 13 center), we are of course less confident.
@@
There are three smaller caveats for Figure 16 also worth noting. First, the confidence intervals assume that the number of local optima at each problem scale are normally distributed. We did not verify as we only saved summary statistics for these experiments. If the local optima count was (for instance) long-tailed toward larger values, this could add uncertainty to our estimates. The second caveat is that we ignore uncertainty in the surrogate solutions’ local optima discovery rate, which is relatively small since the surrogate datasets are orders of magnitude larger than those used to estimate the true discovery rate (2048 and 512 samples for 10 and 20 nodes, respectively). Lastly, 100,000 descents-per-problem may still miss elusive local optima with small basins of attraction. Even so, these local optima are quite unlikely to be global optima if 3-opt trends generalize to higher-dimensional 2-opt solutions [52] (Figure 14 in cited work).
Figure 16: Surrogate solution generation’s estimated mean local optima discovery fraction. Fills show 99% confidence intervals assuming local optima count is normally distributed. True mean local optima count is estimated with small batches of 2-opt local search performing 100,000 descents-per-problem. 10-node surrogate generation never falls below 60% discovery, increasing confidence that best-found local optima are usually the true global optima. 20-node generation falls below 10% discovery by 20 spatial dimensions, suggesting that best-found local optima at these scales may be suboptimal more often.
@@ Appendix G Model architecture
Our model design prioritizes simplicity where possible. Besides making experiments easier to implement, this choice was intended to improve the reproducibility of our results and promote their generalization to other deep architectures. All core model components are implemented using PyTorch [69] Transformer modules. Model code, algorithm code, and training scripts are provided in the project repository at https://github.com/lowellw6/complexity-scaling-laws. Specific hyperparameter choices are discussed separately in Appendix H.
@@ G.1 Policy network
We moved forward with these simplifications after obtaining near-optimal performance without them at our experimented problem scales. But Kool et al.’s design has compelling motivation, especially concerning efficiency. Their graph embedding allows each decoder step to attend over the reduced subset of selectable nodes while still referencing a representation of the full problem. This linear ramp down of attention length still incurs $O(n^{3})$ complexity to decode a full tour, as does our approach which maintains constant attention length. [^9] Self-attention has $O(n^{2})$ complexity w.r.t. to the node scale and attention length $n$ [19], which we iterate $n$ times. But pruning previously selected nodes requires only a third the compute, all else being equal. Given that autoregressive decoding is most of the forward pass, we suggest reconsidering this design choice when experimenting with large node scales.
@@ G.2 Critic network
Attempting to learn more granular credit assignment, we implemented node-level value estimation. We refer to this alternative as compositional value learning, and refer to problem-level value learning as state value learning. We compute compositional values through a straightforward modification to our state-value critic architecture: concatenating the positionally-encoded node selection sequence to the start token for target input. This produces $n+1$ compositional values, each attending over incremental partial tours in one forward pass via a causal mask. The first output remains the problem-level state-value, attending only over the start token and node encodings. The final output is the Q-value, attending over all node selections. We explain how compositional values influence learning objectives in the following subsection.
@@ G.3 PPO using a compositional value baseline
This section describes how we set up the PPO actor-critic loss to learn and leverage node-level compositional value estimates. We use this style of PPO for all RL experiments in the main paper. We omit these details for simplicity and because we observe no evidence of performance differences between node- and problem-level baselines (Figure 17 green and blue).
@@
In theory, this loss still allows compositional value targets to diversify in expectation. If rollouts from a given partial tour usually result in low-cost solutions, the expected compositional value target is small, and the inverse is also true. In preliminary experiments, compositional value learning appeared slightly more sample efficient compared to strictly learning problem-level state values, motivating our selection of the former.
However, this distinction disappears when carefully controlling for other variables and averaging over several seeds, at least for 30-node TSP. After initial transient effects, performance distributions are indistinguishable (Figure 17), and we noticed that compositional value estimates vary little within a given tour. These findings are not surprising given the critic attempts to learn granular values with a coarse reward signal, an analogue to value estimation in a sparsely rewarded MDP. Further, given tours are sampled on-policy for problems generated online, the critic sees each partial tour exactly once.
@@ Appendix H Hyperparameter settings
Setting HPs effectively by hand would be tedious and rely heavily on intuition. Reinforcement learning often involves more HPs than supervised learning, and setting model HPs for our custom architecture lacks the grounding of existing results. Algorithmic HPO has the added benefit of reducing wall-clock time that is required to reach sufficient convergence. With finite resources, unbottlenecked compute can only be approximated in the scaling laws that require it. Some of our intuition-informed preliminary experiments required several million gradient updates (and multiple learning rate warm restarts) to sufficiently converge, dramatically slowing research flow. Our HPO-informed experiments converged well in less than half the gradient updates.
@@ H.1 HPO method
We chose BOHB for its balance of anytime and final performance [72]. BOHB is no longer state-of-the-art in final performance but has comparable anytime performance [74] and is relatively easy to implement. We implemented BOHB using Optuna 3.5’s multivariate Tree-structured Parzen Estimator (TPE) [75] for HP sampling (the Bayesian optimization “BO” in BOHB) paired with their Hyberband [76] pruning algorithm (the “HB” in BOHB). Table 8 details the HPs we used for the BOHB algorithm itself along with our evaluation setup. We found BOHB especially sensitive to the minimum gradient updates budgeted per trial. Setting this value too low results in pruning based on noise before initial convergence. Setting too high wastes a significant amount of time and compute.
@@
Lastly, we found that forcing a warm-start, human-intuition HP configuration for trial 0 allows BOHB to find high-performing regions of the search space considerably faster than when using only random startup trials.